OpenStudy (anonymous):

if 2/3 of the girls in class have brown eyes and 1/4 of the girls have blue eyes, what fraction of the girls in class have neither blue or brown eyes? need help

OpenStudy (anonymous):

2/3 X=R 1/4 X=L remaining=X-2/3 X-1/4 X R and L stands for bRown and bLue

OpenStudy (anonymous):

thanks alot i get now

OpenStudy (anonymous):

Good :)